# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v2

@@
Abstract
Trajectory collection is fundamental for location-based services but often involves sensitive information, such as users’ daily activities, raising significant privacy concerns. Local Differential Privacy (LDP) provides strong privacy guarantees for users, even when the data collector is untrusted. Existing trajectory collection methods under LDP are limited to discrete location spaces, where the number of locations affects both privacy guarantees and trajectory utility. Moreover, many real-world scenarios, such as flying trajectories or sensor trajectories of wearable devices, operate in continuous location spaces, making existing methods inadequate.
This paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. We propose two novel methods: TraCS-D, which perturbs the direction and distance of locations, and TraCS-C, which perturbs the Cartesian coordinates of locations. Both methods are theoretically and experimentally analyzed for trajectory utility in continuous spaces. TraCS can also be applied to discrete spaces by rounding perturbed locations to the nearest discrete points. In this case, TraCS’s privacy and utility guarantees are independent of the number of locations in the space, and has only $\Theta(1)$ time complexity in each perturbation generation. Evaluation results on discrete location spaces validate the efficiency advantage and show that TraCS outperforms state-of-the-art methods with improved trajectory utility, especially for large privacy parameters.
1 Introduction
Trajectory data from users—sequences of locations describing movement over time—are a fundamental resource for activity analysis and location-based services, including activity classification and routine detection. However, directly collecting trajectory data raises significant privacy concerns, especially given users’ growing privacy expectations and stringent regulations [1, 19]. Anonymization techniques like $k$-anonymity [28, 20] have proven insufficient to protect trajectory privacy [6]. Thus, stronger privacy guarantees are necessary for such data.
@@ 1 Introduction
Existing LDP methods for trajectory collection are designed for discrete spaces. They leverage discrete LDP mechanisms, such as the Exponential mechanism [27], to perturb the trajectory data. Discrete LDP mechanisms are explicitly defined on the size of the location space. As a result, these methods either rely on properly splitting the entire location space into grids [29] or assume that the location space consists of a finite set of labeled locations (i.e. points of interest) [9, 37]. However, relying on discrete spaces has three issues: (i) Their privacy guarantees are inherently limited to the chosen discrete sets. For instance, a discrete location space with $10$ points always provides a weaker privacy guarantee than one with $100$ points, since any inference attack has at least a $10\%$ chance of success in the former, regardless of the privacy parameter $\varepsilon$. (ii) Their efficacy and efficiency are often constrained by the size of the discrete set. As the size increases, the probability of selecting the true location decreases. Meanwhile, the widely used Exponential mechanism incurs linear sampling complexity when generating each perturbed location, making it inefficient. (iii) Discrete methods are not directly applicable to naturally continuous location spaces, such as flying and sailing trajectories, sensor trajectories from wearable devices, and similar scenarios.
@@
To address these limitations, this paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. A continuous space models locations as continuous values, such as GPS coordinates in $180,180]\times[-90,90]$. Unlike discrete spaces, it contains infinitely many candidate locations. Collecting trajectory data directly in continuous spaces is natural for many applications, and offers three key advantages over discretizing the space and then applying discrete methods: (i) The privacy guarantee holds for the entire continuous space, regardless of the “number” of locations. (ii) The sampling mechanism operates directly on the continuous space, so its efficacy and efficiency do not relate to the “number” of locations. (iii) Furthermore, perturbed locations in the continuous space can be post-processed by rounding to any discrete space within it, making the approach applicable to both continuous and discrete spaces.
@@
Designing trajectory collection methods for continuous spaces under LDP presents unique challenges: they must ensure LDP for every location in the continuous space; they must preserve sufficient trajectory utility for downstream tasks compared to the original trajectories (formalized in Section 4); and they should efficiently generate perturbed locations to support real-time applications.
@@
This paper proposes two LDP methods for trajectory collection in continuous spaces. The main insight is to decompose the 22D continuous space into 11D subspaces and design mechanisms for each subspace. Our mechanisms leverage the concept of piecewise-based mechanisms [31, 21] for 11D bounded numerical domains. [^1] Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms, e.g. the truncated Laplace mechanism [22, 15]. While they can be incorporated into TraCS, they are more complex and less effective than piecewise-based mechanisms. Section 3.5.4 provides details. Based on two different decompositions of the continuous space, we propose two methods: TraCS-D and TraCS-C. The method TraCS-D treats the continuous space as the composition of a direction space and a distance space. We exploit the direction information of trajectories and design a novel direction perturbation mechanism to ensure LDP for the direction space. Combined with a redesigned piecewise-based mechanism for the distance space, TraCS-D ensures LDP for the continuous 22D space. The method TraCS-C treats the continuous space as the Cartesian product of two distance spaces. Each location is represented by a 22D Cartesian coordinate, and TraCS-C perturbs these coordinates. Both methods guarantee LDP for the entire 22D continuous space. They can also be applied to any discrete space within the continuous space by rounding the perturbed locations to the nearest discrete points. Compared with existing methods based on the Exponential mechanism [9, 37], our methods are unaffected by the number of locations in the location space and require only $\Theta(1)$ time to generate each perturbed location. Specifically, our contributions are as follows:
@@
  • To the best of our knowledge, this is the first work to provide trajectory collection methods for continuous space under pure LDP. We propose TraCS-D and TraCS-C for this purpose. These methods introduce novel location perturbation mechanisms that exploit the direction information and coordinate information of trajectories in continuous spaces. We theoretically and experimentally analyze their trajectory utility.
@@
  • Our methods TraCS can also be applied to discrete spaces. Compared with existing methods for discrete spaces, TraCS has significantly lower computation complexity in generating perturbed locations. Experimental results on discrete location spaces validate the efficiency of TraCS, showing an average time cost of less than $1\%$ of NGram, L-SRR, and ATP [9, 37, 29], three state-of-the-art LDP methods designed for discrete spaces. Additionally, the results demonstrate that TraCS generally achieve better trajectory utility, particularly for larger privacy parameters.
Structure. The rest of the paper is organized as follows. Section 2 introduces the preliminaries. Section 3 presents TraCS-D and TraCS-C methods, their theoretical analysis, along with discussions on their differences and extensions. Following that, Section 4 presents the experimental evaluations on continuous and discrete spaces.
2 Preliminaries
@@
This section formulates the problem, introduces LDP, and reviews three trajectory collection methods under LDP, along with their shortcomings in continuous spaces, which motivates the necessity of designing new LDP methods for trajectory collection in continuous spaces.
2.1 Problem Formulation
@@
2.2 Local Differential Privacy
Definition 1 ($\varepsilon$-LDP [11]).
@@ 2.2 Local Differential Privacy
Theorem 1 (Sequential Composition of LDP [13, 37]).
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two mechanisms that satisfy $\varepsilon_{1}$ and $\varepsilon_{2}$-LDP, respectively. Their composition, defined as $\mathcal{M}_{1,2}:(\mathcal{X}_{1},\mathcal{X}_{2})\to(Range(\mathcal{M}_{1}),Range(\mathcal{M}_{2}))$, satisfies $(\varepsilon_{1}+\varepsilon_{2})$-LDP.
Sequential composition allows us to design LDP mechanisms for 22D data by composing two mechanisms for 11D.
2.2.1 Piecewise-based Mechanism
If $\mathcal{X}$ is continuous and bounded, the state-of-the-art LDP mechanisms are piecewise-based [31, 21], i.e. they sample the output from piecewise probability distributions defined on $\mathcal{X}$. We summarize them in the following form:
@@ 2.2.1 Piecewise-based Mechanism
Piecewise-based mechanisms sample the output $y$ with a high probability $p_{\varepsilon}$ within $l_{x,\varepsilon},r_{x,\varepsilon})$ and a lower probability in the remaining two pieces, while satisfying the $\varepsilon$-LDP constraint. Typical instantiations include PM [31] for mean estimation and SW [21] for distribution estimation. Our perturbation mechanisms also leverage the piecewise-based design but with different instantiations tailored for different data domains.
2.2.2 $k$-$R$ and Exponential Mechanism
If $\mathcal{X}$ is discrete, especially when $|\mathcal{X}|$ is not large [32], a classical LDP mechanism is $k$-$R$ [18].
@@ 2.2.2 $k$-$R$ and Exponential Mechanism
When there is a semantic distance (score) function between elements in $\mathcal{X}$, the Exponential mechanism [27] is more widely used. Unlike $k$-$R$, which treats other values (except $x$) equally, the Exponential mechanism leverages a score function to assign different probabilities. $k$-$R$ is a special case of the Exponential mechanism when the score function is a binary indicator function.
Definition 4 (Exponential Mechanism [27]).
@@
2.3 Existing Methods
Table 1: Comparison with existing methods.
@@ 2.3 Existing Methods
NGram [9] Hierarchical decomposition Discrete space
L-SRR [29] Staircase $R$ mechanism
ATP [37] Direction perturbation
This paper (TraCS)
Direction* &
coordinate perturbation
Continuous space
  • * designed for continuous direction space (different from ATP).
Table 1 summarizes three state-of-the-art trajectory collection methods under pure LDP and our methods.
NGram [9] hierarchically decomposes the physical space $\mathcal{S}$ into fine-grained discrete regions with semantics. Specifically, $\mathcal{S}$ is first decomposed into small regions $R_{s}$, which are further decomposed by categories into $R_{c}$, and then attached with time information $R_{t}$. A decomposed “gram” might look like $r_{sct}=\{\text{mountain, church, 3am}\}$. This decomposition allows trajectory $\mathcal{T}$ to be represented at a knowledge level, enabling the usage of public knowledge to define reachability, i.e. to constrain possible perturbed trajectories. For example, the above gram $r_{sct}$ is unlikely to be visited at 33am as a church is closed at that time, reducing the possible $\mathcal{T}^{\prime}$. Finally, NGram applies the Exponential mechanism to sample a perturbed trajectory from the reachable set.
L-SRR [29] introduces the staircase randomized response (SRR) mechanism for location perturbation. Unlike $k$-$R$, which assigns only two probabilities—one for the true location and another for all other locations—the SRR mechanism employs a staircase-shaped probability distribution. L-SRR recursively partitions the physical space $\mathcal{S}$ and forms location groups $G_{1},G_{2},\dots,G_{g}$ based on their distances to the true location. By integrating the SRR mechanism, L-SRR assigns higher probabilities to locations in groups that are closer to the true location, and lower probabilities to those farther away. This approach is similar to the Exponential mechanism, but uses group-level distance instead of location-level distance.
ATP [37] uses direction perturbation to limit the set of possible locations. It assumes that the location space is a finite set of labeled locations and represents a trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$ as a sequence of these locations. When perturbing a location $\tau_{i+1}$, ATP first perturbs the direction from $\tau_{i}$ to $\tau_{i+1}$. It divides the direction space into $k$ sectors and uses $k$-$R$ to sample a direction sector. Then, the perturbed location $\tau_{i+1}^{\prime}$ is constrained to be within the sampled sector, thereby limiting the possible $\tau_{i+1}^{\prime}$, which is then sampled using the Exponential mechanism.
@@
  • These methods are limited to handling discrete location spaces. The number of locations is a crucial parameter that determines the perturbation mechanism and affects the privacy guarantee. Although it is possible to discretize continuous spaces, proper discretization is challenging to balance the trajectory utility and privacy.
  • Beyond discretization challenges, approaches that rely on the Exponential mechanism (e.g. NGram and ATP) incur high computational cost to generate perturbations and tend to degrade trajectory utility. The other method L-SRR requires forming groups for every location in the space, which leads to the same scalability issue. Additionally, designing effective public knowledge (e.g. NGram) is challenging, particularly in continuous spaces where the location set is infinite and semantic information is difficult to define. Meanwhile, public knowledge is often dataset-specific and also known to the adversary, thereby undermining the privacy guarantee.
Appendix C.1 details the weaknesses of the Exponential mechanism. In summary, before perturbing a location, it requires calculating the score functions $d(x,y)$ of all possible pairs of locations related to $x$. This costs $\Theta(m)$ time, where $m$ is the number of locations in the discrete space. Additionally, each sampling (generating each perturbed location) from the $m$-piece cumulative distribution function (CDF) of the Exponential mechanism also costs $\Theta(m)$ time. This time complexity is not affordable for large-scale location spaces. Moreover, since the sampling probability is determined by the score function, the spatial distribution and relative distances among locations further affect the probability assigned to each location, potentially resulting in utility loss.
To address these limitations, we propose TraCS, a novel trajectory collection method that provides full LDP guarantees for trajectories in continuous spaces. TraCS is specifically designed to overcome the privacy, utility, and efficiency challenges posed by discretization.
3 Our Method: TraCS

### v3

@@
Abstract.
Trajectory collection is essential for location-based services, yet it can reveal highly sensitive information about users, such as daily routines and activities, raising serious privacy concerns. Local Differential Privacy (LDP) offers strong privacy guarantees for users even when the data collector is untrusted. However, existing trajectory collection methods under LDP are largely confined to discrete location spaces, where the size of the location space affects both privacy guarantees and trajectory utility. Moreover, many real-world applications—such as flying trajectories or wearable-sensor traces—naturally operate in continuous spaces, making these discrete-space methods inadequate.
This paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. We propose two methods: TraCS-D, which perturbs the direction and distance of locations, and TraCS-C, which perturbs the Cartesian coordinates of locations. Both methods are theoretically and experimentally analyzed for trajectory utility in continuous spaces. TraCS can also be applied to discrete spaces by rounding perturbed locations to any discrete space embedded in the continuous space. In this case, the privacy and utility guarantees of TraCS are independent of the number of locations in the space, and each perturbation requires only $\Theta(1)$ time complexity. Evaluation results on discrete location spaces validate the efficiency advantage and demonstrate that TraCS outperforms state-of-the-art methods with improved trajectory utility, particularly for large privacy parameters.
local differential privacy, continuous space, trajectory data
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
Trajectory data from users—sequences of locations describing movement over time—are a fundamental resource for activity analysis and location-based services, including activity classification and routine detection. However, directly collecting trajectory data raises significant privacy concerns, especially given users’ growing privacy expectations and stringent regulations (gdpr; DBLP:journals/puc/Krumm09). Anonymization techniques like $k$-anonymity (DBLP:journals/ijufks/Sweene02; DBLP:conf/icde/LiLV07) have proven insufficient to protect trajectory privacy (DBLP:journals/corr/abs-1112-2020). Thus, stronger privacy guarantees are necessary for such data.
@@ 1. Introduction
Existing LDP methods for trajectory collection are designed for discrete spaces. They leverage discrete LDP mechanisms, such as the Exponential mechanism (DBLP:conf/innovations/NissimST12), to perturb the trajectory data. Discrete LDP mechanisms are explicitly defined on the size of the location space. As a result, these methods either rely on properly splitting the entire location space into grids (DBLP:conf/ccs/WangH0QH22) or assume that the location space consists of a finite set of labeled locations (i.e. points of interest) (DBLP:journals/pvldb/CunninghamCFS21; DBLP:journals/pvldb/Zhang [^0] H23). However, relying on discrete location spaces has three inherent limitations: (i) Their privacy guarantees are inherently tied to the discrete set. For example, a space with $10$ locations offers weaker effective privacy than one with $100$ locations: even a trivial inference strategy that always outputs a fixed location succeeds with probability at least $1/10$ in the former, regardless of the privacy parameter $\varepsilon$. [^1] Appendix C.1 provides details on the space-dependence of indistinguishability. (ii) Their efficacy and efficiency are often limited by the domain size. As the number of candidate locations grows, the probability of a mechanism outputting the true location diminishes. Furthermore, the widely used Exponential mechanism incurs linear sampling complexity in the domain size, making the generation of each perturbed location computationally expensive. [^2] Appendix C.2 details the Exponential mechanism’s efficiency and data utility (efficacy) limitations. (iii) Discrete methods are not directly applicable to inherently continuous location spaces, such as those arising from flying and sailing trajectories or sensor data from wearable devices. Although one can discretize a continuous space before applying discrete methods, this inherits the aforementioned limitations. Moreover, choosing a suitable discretization granularity that balances privacy, utility, and efficiency is non-trivial. [^3] Appendix C.3 further discusses the challenges of adapting discrete mechanisms to continuous spaces.
@@
To address these limitations, this paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. A continuous space represents locations as real-valued points, such as GPS coordinates in $180,180]\times[-90,90]$, and thus contains infinitely many candidate locations. Collecting trajectory data directly in continuous spaces is natural for many applications and offers three key advantages. (i) The privacy guarantee does not depend on the cardinality of a discretized location set (ii) The sampling mechanism operates directly on the continuous space, so its efficacy and efficiency are not constrained by the space size. (iii) Perturbed locations can be post-processed (e.g. rounded) to any discrete space embedded within the continuous space, making the approach applicable to both continuous and discrete settings.
@@
Designing LDP trajectory collection methods for continuous spaces poses unique challenges: they must guarantee LDP for every location in the continuous domain; they must preserve sufficient trajectory utility for downstream tasks relative to the original trajectories (formalized in Section 4); and they must generate perturbed locations efficiently to support real-time applications.
@@
This paper proposes two LDP methods for trajectory collection in continuous spaces. The main insight is to decompose the 22D continuous space into 11D subspaces and design mechanisms for each subspace. Our mechanisms build on existing utility-optimized piecewise-based mechanisms (DBLP:journals/popets/ZhengMH25) for 11D bounded numerical domains. [^4] Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms (e.g. the truncated Laplace mechanism (DBLP:journals/tdp/000619; DBLP:journals/jpc/HolohanABA20)). While such mechanisms can be incorporated into TraCS, they are more complex and typically less effective than piecewise-based mechanisms. Section 3.5.6 provides details, and Section 4 includes experimental comparisons. Based on two different decompositions of the continuous space, we propose two methods: TraCS-D and TraCS-C. The method TraCS-D treats the continuous space as the composition of a direction space and a distance space. We exploit the direction information of trajectories and design a direction perturbation mechanism to ensure LDP for the direction space. Combined with a piecewise-based mechanism for the distance space, TraCS-D ensures LDP for the continuous 22D space. The method TraCS-C treats the continuous space as the Cartesian product of two distance spaces. Each location is represented by a 22D Cartesian coordinate, and TraCS-C perturbs these coordinates. Both methods guarantee LDP for the entire 22D continuous space. They can also be applied to any discrete space embedded in the continuous domain by rounding perturbed locations to their nearest discrete points. Compared with existing methods based on the Exponential mechanism (DBLP:journals/pvldb/CunninghamCFS21; DBLP:journals/pvldb/Zhang [^0] H23), TraCS is unaffected by the number of locations in the location space and require only $\Theta(1)$ time to generate each perturbed location. Specifically, our contributions are as follows:
@@
  • To the best of our knowledge, this is the first work to develop trajectory collection methods for continuous spaces under pure LDP. We highlight the benefits of operating directly in continuous spaces (rather than discretizing) and propose TraCS-D and TraCS-C accordingly. Our key insight is to decompose the 22D continuous space into two 11D subspaces and to build 22D trajectory perturbation mechanisms from existing utility-optimized 11D piecewise-based mechanisms, leveraging the direction and coordinate information in continuous trajectories. We theoretically and experimentally analyze their trajectory utility.
@@
  • Our method TraCS also applies to discrete spaces. Compared with existing approaches for discrete spaces, TraCS has substantially lower computational complexity for generating perturbed locations. Experiments on discrete location spaces validate this efficiency, with an average runtime below $1\%$ of that of NGram, L-SRR, and ATP (DBLP:journals/pvldb/CunninghamCFS21; DBLP:journals/pvldb/Zhang [^0] H23; DBLP:conf/ccs/WangH0QH22), three state-of-the-art LDP methods designed for discrete spaces. In addition, the results show that TraCS generally achieves better trajectory utility, particularly at larger privacy parameters.
2. Preliminaries
@@
This section formulates the problem, introduces LDP, and reviews three trajectory collection methods under LDP. We highlight their limitations in continuous spaces, which motivates the need for new LDP methods for trajectory collection in such settings.
2.1. Problem Formulation
@@
2.2. Local Differential Privacy
Definition 0 ($\varepsilon$-LDP (DBLP:conf/focs/DuchiJW13)).
@@ 2.2. Local Differential Privacy
Theorem 2 (Sequential Composition of LDP (DBLP:journals/fttcs/DworkR14; DBLP:journals/pvldb/Zhang [^0] H23)).
Let $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ be two mechanisms that satisfy $\varepsilon_{1}$ and $\varepsilon_{2}$-LDP, respectively. Their composition, defined as $\mathcal{M}_{1,2}\coloneqq(\mathcal{M}_{1},\mathcal{M}_{2}):(\mathcal{X}_{1},\mathcal{X}_{2})\to(Range(\mathcal{M}_{1}),Range(\mathcal{M}_{2}))$, satisfies $(\varepsilon_{1}+\varepsilon_{2})$-LDP.
Sequential composition enables (i) designing LDP mechanisms for 2D data by composing two 1D mechanisms, and (ii) perturbing multiple locations in a trajectory (i.e. repeated composition) while still providing an overall LDP guarantee.
2.2.1. Piecewise-based Mechanism
When the input domain $\mathcal{X}$ is continuous and bounded, state-of-the-art LDP mechanisms are often piecewise-based (DBLP:conf/icde/WangXYZHSS019; DBLP:conf/sigmod/Li0LLS20; DBLP:journals/popets/ZhengMH25). They generate an output by sampling from a probability distribution over $\mathcal{X}$ whose density is defined in a piecewise manner. Below, we present a unified formulation that captures this family of mechanisms.
@@ 2.2.1. Piecewise-based Mechanism
Piecewise-based mechanisms sample an output $y$ with a high probability $p_{\varepsilon}$ from the interval $l_{x,\varepsilon},r_{x,\varepsilon})$, and with a lower probability from the remaining two pieces, while still satisfying $\varepsilon$-LDP. Representative instantiations include OGPM (DBLP:journals/popets/ZhengMH25) and PM (DBLP:conf/icde/WangXYZHSS019) for mean estimation, and SW (DBLP:conf/sigmod/Li0LLS20) for distribution estimation. Our perturbation mechanisms build on the 1D piecewise-based mechanism in OGPM to design new 2D mechanisms tailored to trajectory collection.
2.2.2. $k$-$R$ and Exponential Mechanism
If $\mathcal{X}$ is discrete, especially when $|\mathcal{X}|$ is not large (DBLP:conf/uss/WangBLJ17), a classical LDP mechanism is $k$-$R$ (or GRR in some literature) (DBLP:conf/nips/KairouzOV14).
@@ 2.2.2. $k$-$R$ and Exponential Mechanism
When there is a semantic distance (score) function between elements in $\mathcal{X}$, the Exponential mechanism (DBLP:conf/innovations/NissimST12) is more widely used. Unlike $k$-$R$, which treats other values (except $x$) equally, the Exponential mechanism leverages a score function to assign different probabilities. $k$-$R$ is a special case of the Exponential mechanism when the score function is a binary indicator function.
Definition 0 (Exponential Mechanism (DBLP:conf/innovations/NissimST12)).
@@
2.2.3. Privacy Parameter for Trajectory Data
When applying LDP to trajectory data, there are two strategies for setting the privacy parameter. (i) Treating each location $\tau_{i}$ in a trajectory $\mathcal{T}$ as a data item and applying an $\varepsilon$-LDP mechanism to perturb each location; (ii) Treating the entire trajectory $\mathcal{T}$ as a data item and applying an $\varepsilon$-LDP mechanism to perturb the whole trajectory. [^5] This concerns how to set the privacy parameter for trajectory data, rather than defining the “trajectory space” as the input space in LDP mechanisms. We focus on LDP mechanisms for location spaces in this paper, as the trajectory space in $\mathbb{R}^{2}$ is infinite-dimensional. The first strategy is cleaner and convertible to the second strategy via the Sequential Composition Theorem 2, whereas the second depends on the trajectory length, which complicates the analysis and the design of mechanisms under a fixed $\varepsilon$. Therefore, we primarily adopt the first strategy in this paper, and include experiments for the second in Appendix C.10 for completeness.
2.3. Existing Methods
Table 1. Comparison with existing methods.
@@ 2.3. Existing Methods
NGram (DBLP:journals/pvldb/CunninghamCFS21) Hierarchical decomposition Discrete space
L-SRR (DBLP:conf/ccs/WangH0QH22) Staircase RR mechanism
ATP (DBLP:journals/pvldb/Zhang [^0] H23) Direction perturbation
This paper (TraCS) Directiona  & coordinate perturbation Continuous space
  • a
    Designed for continuous direction space (different from ATP).
Table 1 summarizes three state-of-the-art trajectory collection methods under pure LDP, along with our approach. Other works that emphasize external knowledge rather than new perturbation mechanisms are discussed in Appendix A.
NGram (DBLP:journals/pvldb/CunninghamCFS21) hierarchically decomposes the physical space $\mathcal{S}$ into fine-grained discrete regions with semantic labels. Concretely, $\mathcal{S}$ is first partitioned into small spatial regions $R_{s}$, then further refined into category-specific regions $R_{c}$, and finally augmented with temporal information $R_{t}$. A resulting “gram” can be written as $r_{sct}=\{\text{mountain, church, 3am}\}$. This decomposition represents a trajectory $\mathcal{T}$ at the knowledge level, enabling the use of public knowledge to define reachability, i.e. to constrain the set of feasible perturbed trajectories. For instance, visiting a church at 33am is unlikely because it is typically closed, which reduces the reachable set of $\mathcal{T}^{\prime}$. Finally, NGram applies the Exponential mechanism to sample a perturbed trajectory from this reachable set.
L-SRR (DBLP:conf/ccs/WangH0QH22) introduces the staircase randomized response (SRR) mechanism for location perturbation. Unlike $k$-$R$, which assigns only two probabilities—one for the true location and another for all other locations—the SRR mechanism employs a staircase-shaped probability distribution. L-SRR recursively partitions the physical space $\mathcal{S}$ and forms location groups $G_{1},G_{2},\dots,G_{g}$ based on their distances to the true location. By integrating the SRR mechanism, L-SRR assigns higher probabilities to locations in groups that are closer to the true location, and lower probabilities to those farther away. This approach is similar to the Exponential mechanism, but uses group-level distance instead of location-level distance.
ATP (DBLP:journals/pvldb/Zhang [^0] H23) perturbs directions to restrict the set of feasible perturbed locations. For a trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, when perturbing $\tau_{i+1}$, ATP first perturbs the direction from $\tau_{i}$ to $\tau_{i+1}$. Specifically, it partitions the direction space into $k$ sectors and applies $k$-$R$ to sample a sector. The perturbed location $\tau_{i+1}^{\prime}$ is then constrained to lie within the sampled sector, which reduces the candidate set of $\tau_{i+1}^{\prime}$; ATP finally samples $\tau_{i+1}^{\prime}$ from this constrained set using the Exponential mechanism.
@@
  • Existing methods are designed for discrete location spaces. Discretizing a continuous domain is possible, but choosing an appropriate granularity is non-trivial.
  • Beyond discretization, methods that rely on the Exponential mechanism (e.g. NGram and ATP) typically incur high computational cost for evaluating scores and sampling, and their utility can be sensitive to the spatial distribution of candidate locations. L-SRR requires constructing distance-based groups for every location, leading to similar scalability issues. Moreover, approaches that depend on public knowledge (e.g. NGram) are hard to design in continuous spaces (where the location set is infinite and semantics are difficult to define), are often dataset-specific, and the public knowledge is typically known to the adversary, which can further weaken privacy protection.
Appendix C.2 details the limitations of the Exponential mechanism. In brief, perturbing a location requires evaluating the score function $d(x,y)$ for all candidate outputs $y$ associated with $x$, which takes $\Theta(m)$ time where $m$ is the number of locations in the discrete space. Moreover, generating each perturbed location requires sampling from an $m$-piece cumulative distribution function (CDF), which also incurs $\Theta(m)$ time. Such a per-location cost is prohibitive for large-scale location spaces. Finally, because the sampling probabilities are induced by the score function, they can be sensitive to the spatial distribution of candidate locations and their relative distances, which may further degrade utility.
To address these limitations, we propose TraCS, a novel trajectory collection method that ensures rigorous LDP guarantees for trajectories in continuous spaces. TraCS is carefully designed to tackle the privacy, utility, and efficiency challenges inherent in handling continuous domains without relying on discretization.
3. Our Method: TraCS